# A Search for Classical Subsystems in Quantum Worlds

### v2

@@ IV Results
We numerically searched for factorizations using a number of different schemes. Each scheme draws from one of three approaches to initial conditions, and one of four types of Hamiltonian. In this section we present these technical findings. In what follows, it is important to remember that there are two factorizations of the Hilbert space that one must keep track of: the “fiducial” factorization, $\mathcal{H_{\sigma}}\otimes\mathcal{H_{\epsilon}}$ in which the Hamiltonian is first supplied, and the “destination” $\mathcal{H}_{s}\otimes\mathcal{H}_{e}$ which is the one our algorithm lands on. The two are related by the unitary $B$ as in Eq. (4).
IV.1 Approaches to initial conditions
While we are interested in finding the optimal factorization, parameterized by the unitary $B$, our cost, Eq. (10), is a function of both $B$ and the initial state $|\psi(0)\rangle_{w}$. Thus, while we always optimize over $B$, we explore various levels of optimizing the initial state which we outline below in increasing order of complexity.
No initial state optimization,
Here we fix our initial state to be a product state in the fiducial basis $|\psi(0)\rangle_{w}=|\xi\rangle_{\sigma}|\phi\rangle_{\epsilon}$ and optimize over just the factorization parameterized by $B$. This case corresponds to asking whether a set of dynamics, with a particular initial condition, can be viewed as corresponding to two approximately well-defined subsystems.
System state optimization.
In addition to optimizing over $B$, we also optimize over the initial state in the fiducial system Hilbert space, $\mathcal{H}_{\sigma}$, i.e. $|\psi(0)\rangle_{w}=A_{\sigma}\otimes 1_{\epsilon}(|\xi\rangle_{\sigma}|\phi\rangle_{\epsilon})$ where $A_{\sigma}\in U(d_{s})$ is to be optimized. This is typically the case encountered in the lab: the experimentalist can set the state of the system being studied, for example the spin of an electron, and not of the more disordered environment with which it interacts.
However, in the case of cosmology, it is not clear what such a bipartite factorization should look like. For example, in order to study the decoherence of the curvature perturbations that are setup during the inflationary epoch, some have defined the modes that leave the Hubble horizon (i.e. the ones that leave observable imprints on large-scale structure) as the “system” while modes that stay within the Hubble horizon as the “environment”, though this choice remains arbitrary [26, 27, 28, 29, 30]. [^6] See also [23] where it is argued that it is not necessary to specify an “environment” in order to explain the emergent classicality for inflationary perturbations (due to quantum squeezing [21]). Or as the authors call it “decoherence without decoherence”. One approach might be to make choices that reflect only quantities within the limits of what we actually measure, but the Universe as we usually model it behaves very classically even well outside those limits. Still, those limits are steadily receding in the face of new astronomical observations and new quantum phenomena, such as non-gaussianity or entanglement [31], could emerge as the new data comes in.
Global state optimization.
In this approach, in addition to optimizing over the factorization parameterized by $B$, we optimize the global initial state, i.e $|\psi(0)\rangle_{w}=A_{w}|\xi\rangle_{\sigma}|\phi\rangle_{\epsilon}$ where $A_{w}\in U(d_{w})$. By allowing ourselves freedom in the global initial state we can stay oblivious to any prior notion of a working basis and, in cases where a notion of a system-environment factorization is not defined a priori, allow us to define the factorization in which a random initial system state decoheres.
IV.2 Hamiltonians considered
The existence of a subsytem description depends on the dynamics governing the evolution of the global state. The discussion of the sort of dynamics that enable the existence of pointer states is deferred to Section IV.4; here we merely state our trial Hamiltonians. Note that a particular Hamiltonian is fundamentally characterized by its eigenvalues and any physical description is attributed to the choice of a fiducial factorization.
@@ IV.2 Hamiltonians considered
Here we imagine that there is a central qubit in $\mathcal{H_{\sigma}}$ that interacts with a bath of environment qubits in $\mathcal{H_{\epsilon}}$, where the interaction with each environment qubit consists of a single Pauli string, in the presence of an external field. That is,
@@
where $\sigma^{(.)}\in\{\sigma^{X},\sigma^{Y},\sigma^{Z}\}$ is a single qubit Pauli operator and we have suppressed the identity operators in each term. The central spin model is of interest to us because its parameters can be easily dialed to study various cases that appear in the decoherence literature. For example, setting $\beta_{i}=0$ gives a perfectly decoupled Hamiltonian or, by choosing each $p_{i}$ from $\{X,Y,Z\}$ such that $p_{i}\neq p_{j}$ (for $n_{e}\leq 3$) one can have a simple Hamiltonian which nonetheless contains several non-commuting terms. We ran both of these sub-categories through our algorithm.
Quantum Measurement Limit
This Hamiltonian takes the form $H=H_{\sigma}\otimes H_{\epsilon}$ in the fiducial factorization where $H_{\sigma/\epsilon}$ are random (traceless) Hermitian matrices. We use the term “quantum measurement limit” here, although the tensor product form is actually a (commonly studied) special case of this limit (which is discussed in more generality in [32])
@@
IV.3 Results for each of the twelve combinations of the above
We ran the four Hamiltonians in Sec. IV.2, for each of the three approaches to initial conditions in Sec. IV.1, $\mathcal{O}(100)$ times to explore the solution space obtained by our algorithm. For each iteration, we generate a starting guess randomly for the unitaries $B$ and $A$ (where applicable), as well as generating a new Hamiltonian for the random Hamiltonian and quantum measurement limit cases. We summarize our results in the tile plot of Fig. 2 where we select the run achieving the lowest cost in the late time limit for each of the categories. We take this late-time limit to be $t\approx 10^{5.5}$ due to numerical considerations (as discussed in Appendix C). Let us understand these statistics row-by-row.
The top row of Fig. 2 shows results for the case where there is no optimization over the unitary $A$ (defined only in the system and global state optimization cases). For that case the algorithm can find a satisfactorily low cost for only the decoupled Hamiltonian where any unitary of the form $B=B_{\sigma}\otimes B_{\epsilon}$ will form a solution (although this does not imply that all solutions to $B$ must be tensor products; consider a SWAP gate for a trivial counter-example). Unsurprisingly, we do not get low cost solutions for the other Hamiltonians with this initialization approach; even though the cost is $\mathcal{O}(10^{-4})$ for the non-commuting central spin (second column), the solution does not generalize to arbitrary times (i.e. other than the ones on which it was trained). While not shown here, this non-generalizability can be seen in Fig. 7 in Appendix C.
@@ IV.3 Results for each of the twelve combinations of the above
Finally, the last row shows results for the case when we optimize over the global initial state. In this case, the algorithm is successful in finding generalizable solutions to all categories of Hamiltonians. There is a wide variety of solutions, each with a different physics interpretation, that are available to the algorithm in this scenario; we elaborate more on the details of this solution space in Sec. IV.4.
While Fig. 2 only shows the most optimal solution, the detailed statistics for each class of Hamiltonians is discussed in Appendix C and Fig. 7 therein.
IV.4 Analysis
Figure 3: Time evolution of the purity of the reduced system density matrix for each of the categories of the “destination” Hamiltonians identified by our algorithm, as discussed in Sec. IV.4. When the purity is close to one, the training state is separable, i.e. $|\psi\rangle_{w}=|\chi\rangle_{s}|\phi_{e}\rangle$ in the destination factorization. Thus, we also show the effect on the purity resulting from substituting the system state for the orthogonal state, $|\chi\rangle_{s}\rightarrow|\chi^{\prime}\rangle_{s}$, and of randomizing the environment state $|\phi\rangle_{e}$. While in the decoupled case all initial product states remain unentangled for arbitrary times, the other destinations only allow very particular states to stay unentangled beyond the decoherence time. The furnace case is peculiar in that it admits a pointer state only for particular choices of the environment state, as elaborated on in Fig. 5.
There are a plethora of solutions that can satisfy the problem of identifying subsystems as we have posed it here (i.e. by the vanishing of the cost Eq. (9)). Below, we outline these categories and elaborate on the characteristics of each of the class of solutions. We are particularly interested in the dynamics that enable the existence of pointer states and here we adopt the “Heisenberg picture” where we study the Hamiltonian in the new factorization. The properties exhibited by these various “destination” Hamiltonians can also be inferred from the evolution of different kinds of initial states as shown in Fig. 3. In this section we discuss each type of destination Hamiltonian in turn, enumerated in order of the corresponding panel in Fig. 3
@@ IV.4 Analysis
so that the system and environment are totally decoupled. In this case, the entropy of the initial state remains constant and an initial product state remains unentangled. However, this is a vanishingly rare solution that is not expected for random Hamiltonians. This can be understood heuristically by a simple degrees-of-freedom counting argument: decomposing a generic Hamiltonian in the Pauli basis gives a total of $4^{n_{w}}$ terms (where $n_{w}$ is the total number of qubits) which can be arranged in a matrix of Pauli coefficients of dimensions ($4^{n_{s}}$, $4^{n_{e}}$) in such a way that the first column and row correspond to the system and environment self-interaction terms respectively. Thus, there are a total of $14^{n_{s}}-$ 1 terms that contribute to the system self-interaction and $14^{n_{e}}-$ 1 for the environment (the $-1$ occurs because there is a Pauli coefficient for the identity which only contributes a global phase and is not of physical significance). The rest of the $14^{n_{s}+n_{e}}-4^{n_{s}}-4^{n_{e}}+$ 1 terms contribute to the interaction Hamiltonian, $H_{\rm int}$. Thus, say for a 3-qubit world, expecting to find a factorization where the Hamiltonian is decoupled amounts to the expectation that the $45$ interaction coefficients can be regrouped into $20$ slots for the self-interaction terms 777 Another way the decoupled form appears difficult to realize is that in the decoupled case the eigenvalues of $H_{w}$ are sums of pairs of the $n_{s}$ eigenvalues of $H_{s}$ and the $n_{e}$ eigenvalues of $H_{e}$. This creates rigid constraints on the $n_{w}$ eigenvalues of $H_{w}$, and it is not clear that such constraints can be generally realized. See [33] for a related investigation..
@@
It turns out that the algorithm finds this solution for the (interacting) central spin model. This may initially seem surprising, however, in fact this model passes the counting argument since, in the fiducial factorization, $H_{\rm int}$ contains only $n_{e}$ non-zero Pauli terms. We depict the result of evolving several product states in this decoupled factorization of the central spin model in Fig. 3; clearly they all remain product states throughout the evolution.
@@
The transformation of an interacting Hamiltonian to a decoupled one is familiar from bosonization of interacting field theories [34, 35] or the Jordan-Wigner (JW) transformation. The latter is particularly pertinent for our central spin Hamiltonian since, for several classes of Hamiltonians that are quadratic in the spin operators, the JW transformation unitarily transforms the spin system to a set of free fermions [36, 37]. We comment on the interpretation of such non-local transformations in Section V. Some authors have argued that factorizations with quasi-classical subsystems can be identified as those that minimize the interaction Hamiltonian [3, 7]. While this decoupled case certainly passes that criterion, the requirement seems too restrictive. In the rest of this section, we show several destination Hamiltonians that admit pointer states, and thus a notion of quasi-classicality, despite having significant interactions between the subsystems. We refer the reader to Ref. [38] for an implementation of a quantum algorithm that can be used to search for factorizations in which the Hamiltonian is decoupled.
@@
For any Hamiltonian, there always exists a factorization, defined by the eigenstates of the Hamiltonian, that admits pointer states. Specifically, we show in Appendix A that for any Hamiltonian there always exists a factorization of the Hilbert space where the eigenstates are separable so that it may be written in the form
@@
It is straightforward to verify, as shown in Appendix A, that $|\chi_{i}\rangle_{s}$ are the pointer states and that a generic state will evolve into a state whose system density matrix is diagonal in the pointer basis, i.e.
@@
(for a sufficiently large and scrambling environment) as is characteristic of einselection. It is also manifest from the form of Eq. (14) that $H$ is block diagonal in the pointer basis and, furthermore, that the $|\chi_{i}\rangle_{s}$ are robust to entanglement independent of the choice of the environment state or the size or scrambling nature of the environment. [^8] The distinction we are making here is that the robustness addresses the evolution of a state where the system is initially in the pure state $|\chi_{i}\rangle_{s}$, and Eq. (15) reflects the behavior of an arbitrary initial state. This feature can be seen in Fig. 3 where we show that randomizing the environment state does not degrade the purity of the reduced density matrix. We also note that the pointer states are stationary, while an arbitrary environment state will evolve under the evolution generated by the $H_{e}^{(i)}$ operators. All these properties, as well as a general recipe for finding this factorization for an arbitrary Hamiltonian, are proved in Appendix A.
Such a mapping to a factorization where the energy eigenstates are separable has some similarity with the case of dynamical decoupling in superconducting qubits (see e.g. Appendix A in [39] for an overview) . There, the “logical” qubits (what we call the “destination” TPS) correspond to the energy eigenstates of a world described by interacting “physical” qubits (what we have called the “fiducial” TPS). Defining the logical qubits in this way can mitigate certain sources of noise[40].
In the decoherence literature authors often consider special Hamiltonians; such as those for which the system self-interaction commutes with the interaction, $[H_{s}\otimes 1_{e},H_{\rm int}]=0$ or a “quantum measurement limit” where the Hamiltonian is dominated by a single interaction term $H\approx H_{s}\otimes H_{e}$. These are both special cases of the more general block-diagonalization procedure outlined above where the $H_{e}^{(j)}$ operators in Eq. (14) have specific relationships. For example, in the quantum measurement limit all the $H_{e}^{(i)}$ operators are the same up to an overall scaling, $H_{e}^{(j)}\propto H_{e}^{(i)}$. A particularly interesting sub-case of the block diagonal Hamiltonian is when there exists a decoherence-free subspace (DFS) (a topic relevant for quantum computing [41, 42]). In this factorization some, but not all, eigenvectors have vanishing entropy and, as in the quantum measurement limit, their corresponding $H_{e}^{(i)}$ operators are proportional to each other. Writing out the Hamiltonian in the global eigenbasis, we can interpret it as a sum of a decoupled (i.e. DFS) and an entangling part:
@@
Such a DFS solution lies in the intermediate regime of the general block diagonal form in Eq. (14) and the decoupled Hamiltonian in Eq. (13). Since $H_{\rm DFS}$ is subject to the same constraints as those for the (fully) decoupled Hamiltonian (see footnote 7), we do not expect to find such a factorization for random Hamiltonians. However, the central spin model admits both the decoupled and the block diagonal destinations, and in that case our algorithm is indeed able to find this intermediate DFS solution.
@@
This class of solutions is particularly interesting because it exhibits dynamic pointer states, though it fails to provide a sufficiently distinguishable basis of such states. Here, there is a dominant contribution to $|\psi\rangle_{w}$ from a low entropy energy eigenstate and smaller but, crucially, non-negligible contributions from high entropy eigenstates. More precisely, in the language of Appendix A, the world energy eigenstates are of the form,
@@
where the Schmidt coefficients $\beta_{j}^{(i)}$ are not all zero for $j>1$ $\forall i\in\{2,..,d_{w}\}$ (for convenience we choose to label the low entropy eigenstate $|1\rangle_{w}$). And the global state is of the form,
@@
This decomposition of $|\psi\rangle_{w}$ into eigenstates $|i\rangle_{w}$ and their corresponding linear entropy of entanglement, Eq. (6), is shown in Fig. 4 for a world with one system and two environment qubits. We note that the correlation between the entropy and $|\alpha_{i}|^{2}$ for $i>2$ seems typical of these ‘approximate eigenstate’ solutions but that it is likely a feature of the search algorithm and not of inherent physical significance. We have checked numerically that, as long as the condition $\alpha_{1}\gg\alpha_{j>1}$ is satisfied, there is considerable freedom in distributing the weights amongst the other higher entropy eigenstates. In Fig. 3 c. we show the evolution of the purity of the redcued system density matrix of such a state (labeled ‘training $|\psi\rangle_{w}$’) compared to other states.
Figure 4: The decomposition of $|\psi\rangle_{w}$ in the energy eigenbasis for the case of the ‘approximate eigenstate” solutions discussed in Sec. IV.4 (see Eq. (19)) is shown in blue along with their corresponding linear entropy of each energy eigenstate (defined in Eq. (6)). These solutions are dominated by a single low entropy energy eigenstate, although high entropy eigenstates contribute to a small degree. The plot is arranged in order of decreasing $|\alpha_{i}|$.
The fact that the contribution to $|\psi\rangle_{w}$ from the high entropy eigenstates is small, but not negligible, is crucial as it leads to a pure $\rho_{s}$ for times several orders of magnitude greater than the decoherence time (see Fig. 3 c.), while allowing the state to be dynamic. These two behaviors resonate with our classical experience of subsystems. It appears, at first, that such solutions are exactly the ones that the author of [3] is in search of, i.e. those that have a high degree of predictability (as evidenced by the purity of the system state) as well as “autonomy” (that their dynamics are not trivial). However, while this class of solutions satisfactorily minimizes our cost function (Eq. (10)), it is not clear how to apply the decoherence framework to it since we do not have a multi-dimensional distinguishable basis of states of the form in Eq. (19). This should be contrasted with the TPS that leads to a block diagonal Hamiltonian in the pointer basis (c.f. part (b)) where we do have a complete pointer basis (such that the system density matrix follows Eq. (15)) but that the pointer states remain static.

### v3

@@ IV Results
We now turn to our main findings from the computational and analytic explorations. We focus on four types of Hamiltonian and, for the numerical work, use all three initialization approaches delineated in Sec. III.2 for each Hamiltonian. Sec. IV.2 summarizes results from the algorithm (with detailed statistics deferred to Appendix D) and Sec. IV.3 identifies these numerical results with analytic forms that the Hamiltonian takes in the new TPS.
In what follows, it is important to remember that there are two factorizations of the Hilbert space that one must keep track of: the “fiducial” factorization, $\mathcal{H_{\sigma}}\otimes\mathcal{H_{\epsilon}}$ in which the Hamiltonian is first supplied, and the “destination” $\mathcal{H}_{s}\otimes\mathcal{H}_{e}$ which is the one we land on. The two are related by the unitary $B$ as in Eq. (4).
IV.1 Hamiltonians considered
The existence of a subsytem description depends on the dynamics governing the evolution of the global state. The discussion of the sort of dynamics that enable the existence of pointer states is deferred to Section IV.3; here we merely state our trial Hamiltonians. Note that a particular Hamiltonian is fundamentally characterized by its eigenvalues and any physical description is attributed to the choice of a fiducial factorization.
@@ IV.1 Hamiltonians considered
Here we study a system with a central qubit in $\mathcal{H_{\sigma}}$ that interacts with a bath of environment qubits in $\mathcal{H_{\epsilon}}$, where the interaction with each environment qubit consists of a single Pauli string, in the presence of an external field. That is,
@@
where $\sigma^{(.)}\in\{\sigma^{X},\sigma^{Y},\sigma^{Z}\}$ is a single qubit Pauli operator and we have suppressed the identity operators in each term. The central spin model is of interest to us because its parameters can be easily modified to study various cases that appear in the decoherence literature. For example, setting $\beta_{i}=0$ gives a perfectly decoupled Hamiltonian or, by choosing each $p_{i}$ from $\{X,Y,Z\}$ such that $p_{i}\neq p_{j}$ (for $n_{e}\leq 3$) one can have a simple Hamiltonian which nonetheless contains several non-commuting terms. We studied both of these sub-categories with our algorithm.
Quantum Measurement Limit.
This Hamiltonian takes the form $H=H_{\sigma}\otimes H_{\epsilon}$ in the fiducial factorization where $H_{\sigma/\epsilon}$ are random (traceless) Hermitian matrices.  [^7] We use the term “quantum measurement limit” here, although the tensor product form is a special case of this limit (which is discussed in more generality in [schlosshauer2007decoherence]).
@@
IV.2 Numerical results for the twelve combinations of Hamiltonians & initial conditions
Figure 1: Best late time linear entropy achieved by the different optimization methods for various Hamiltonians. The columns correspond to the Hamiltonians described in Sec. IV.1. The rows correspond to the different initial state optimizations from Sec. III.2; the unitary $B$ is optimized in each case.
We ran the four Hamiltonians in Sec. IV.1, for each of the three approaches to initial conditions in Sec. III.2, $\mathcal{O}(100)$ times to explore the solution space obtained by our algorithm. For each iteration, we generate a starting guess randomly for the unitaries $B$ (parameterizing the TPS) and $A$ (parameterizing the initial state; as needed), as well as generating a new Hamiltonian for the random Hamiltonian and quantum measurement limit cases. We summarize our results in the tile plot of Fig. 1 where we select the run achieving the lowest cost in the late time limit for each of the categories. We take this late-time limit to be $t\approx 10^{5.5}$ due to numerical considerations (as discussed in Appendix D). Let us understand these statistics row-by-row.
The top row of Fig. 1 shows results for the case where there is no optimization over the unitary $A$ (defined only in the system and global state optimization cases). For that case the algorithm can find a satisfactorily low cost for only the decoupled Hamiltonian where any unitary of the form $B=B_{\sigma}\otimes B_{\epsilon}$ will form a solution (although this does not imply that all solutions to $B$ must be tensor products; consider a SWAP gate for a trivial counter-example). Unsurprisingly, we do not get low cost solutions for the other Hamiltonians with this initialization approach; even though the cost is $\mathcal{O}(10^{-3})$ for the non-commuting central spin (second column), the solution does not generalize to arbitrary times (i.e. other than the ones on which it was trained). This non-generalizability can be seen in Fig. 7 of Appendix D.
@@ IV.2 Numerical results for the twelve combinations of Hamiltonians & initial conditions
Finally, the last row shows results for the case when we optimize over the global initial state. In this case, the algorithm is successful in finding generalizable solutions to all categories of Hamiltonians. There is a wide variety of solutions, each with a different physics interpretation, that are available to the algorithm in this scenario; we elaborate more on the details of this solution space in Sec. IV.3.
While Fig. 1 only shows the most optimal solution, the detailed statistics for each class of Hamiltonian is discussed in Appendix D and Fig. 7 therein. The numerical results presented here are shown for a single qubit system and a two qubit environment ($n_{s}=1$, $n_{e}=2$, an $8$ dimensional Hilbert space). We explored a variety of sizes up to $n_{s}=2$ and $n_{e}=4$ (a $64$ dimensional Hilbert space) and saw similar results. In the next section we demonstrate, using analytic forms of the destination Hamiltonian, how the low costs described here can be achieved by the algorithm. In all but one case, the analytic forms exist for arbitrary $n_{s}$ and $n_{e}$.
IV.3 Analysis
Figure 2: Time evolution of the purity of the reduced system density matrix for each of the categories identified by our algorithm, as discussed in Sec. IV.3. When the purity is close to one, the training state is separable, i.e. $|\psi\rangle_{w}=|\chi\rangle_{s}|\phi_{e}\rangle$ in the destination factorization. Thus, we also show the effect on the purity resulting from substituting the system state for the orthogonal state, $|\chi\rangle_{s}\rightarrow|\chi^{\prime}\rangle_{s}$, and of randomizing the environment state $|\phi\rangle_{e}$. While in the decoupled case all initial product states remain unentangled for arbitrary times, the other destinations only allow very particular states to stay unentangled beyond the decoherence time. The furnace case is interesting in that it admits a pointer state only for particular choices of the environment state, as elaborated on in Fig. 4.
There are many solutions that can satisfy the problem of identifying subsystems as we have posed it here by the vanishing of the cost given in Eq. (12). In this section we offer a physical understanding of the different solutions we have found. We adopt the “Heisenberg picture” where we study the Hamiltonian in the new factorization.
We have found the solutions can be grouped in five categories, each with its own characteristic behaviors. The Decoupled Hamiltonian category is especially simple, but these solutions only exist for special cases. The Block Diagonal Hamiltonian solutions exist even for completely random fiducial Hamiltonians, and correspond to perfectly stable static pointer states. The Approximate Eigenstate, Furnace Hamiltonian and Extended coherence time each have more nuanced dynamics, and our understanding of these categories is a bit more phenomenological.
Next, we discuss each category in turn, elaborating on the characteristics these solutions exhibit and also comment on their scalability to larger systems. Figure 2. shows the behaviors of specific solutions from each category, and is used to support the discussion which follows. Which category of solution is found on a specific run of our algorithm involves an element of chance, as well as certain constraints which we also delineate below.
@@ IV.3 Analysis
so that the system and environment are totally decoupled. In this case, the entropy of the initial state remains constant and an initial product state remains unentangled. However, this is a vanishingly rare solution that is not expected for random Hamiltonians. This can be understood heuristically by a simple degrees-of-freedom counting argument: decomposing a generic Hamiltonian in the Pauli basis gives a total of $4^{n_{w}}$ terms (where $n_{w}$ is the total number of qubits) which can be arranged in a matrix of Pauli coefficients of dimensions ($4^{n_{s}}$, $4^{n_{e}}$) in such a way that the first column and row correspond to the system and environment self-interaction terms respectively. Thus, there are a total of $4^{n_{s}}-1$ terms that contribute to the system self-interaction and $4^{n_{e}}-1$ for the environment (the 1-1 occurs because there is a Pauli coefficient for the identity which only contributes a global phase and is not of physical significance). The rest of the $4^{n_{s}+n_{e}}-4^{n_{s}}-4^{n_{e}}+1$ terms contribute to the interaction Hamiltonian, $H_{\rm int}$. Thus, say for a 3-qubit world, expecting to find a factorization where the Hamiltonian is decoupled amounts to the expectation that the $45$ interaction coefficients can be regrouped into $20$ slots for the self-interaction terms  [^8] Another way the decoupled form appears difficult to realize is that in the decoupled case the eigenvalues of $H_{w}$ are sums of pairs of the $n_{s}$ eigenvalues of $H_{s}$ and the $n_{e}$ eigenvalues of $H_{e}$. This creates rigid constraints on the $n_{w}$ eigenvalues of $H_{w}$, and it is not clear that such constraints can be generally realized. See [knutson2000honeycombs] for a related investigation..
@@
A TPS that realizes a decoupled Hamiltonian cannot generally be found when starting with a random Hamiltonian, but our algorithm does find such a TPS for the interacting central spin model. This case does pass the counting argument test since in the fiducial factorization, $H_{\rm int}$ contains only $n_{e}$ non-zero Pauli terms. We depict the result of evolving several product states in this decoupled factorization of the central spin model in Fig. 2; clearly they all remain product states throughout the evolution.
@@
The transformation of an interacting Hamiltonian to a decoupled one is familiar from bosonization of interacting field theories [coleman1975quantum-bosonization, von1998bosonization-reformionization] or the Jordan-Wigner (JW) transformation. The latter is particularly pertinent for our central spin Hamiltonian since, for several classes of Hamiltonians that are quadratic in the spin operators, the JW transformation unitarily transforms the spin system to a set of free fermions [lieb1961two-XY_JW, schultz1964two-lieb-Ising_JW]. We comment on the interpretation of such non-local transformations in Section V. Some authors have argued that factorizations with quasi-classical subsystems can be identified as those that minimize the interaction Hamiltonian [tegmark2015consciousness, zanardi2022operational]. Ref. [mansuroglu2024quantum] presents a quantum algorithm optimized to search explicitly for factorizations in which the interaction term is minimized. While the decoupled case certainly passes that criterion, the requirement seems too restrictive. In the rest of this section, we show several destination Hamiltonians that admit pointer states, and thus a notion of quasi-classicality, despite having significant interactions between the subsystems.
@@
For any Hamiltonian, there always exists a factorization, defined by the eigenstates of the Hamiltonian, that admits pointer states. Specifically, we show in Appendix B that for any Hamiltonian there always exists a factorization of the Hilbert space where the eigenstates are separable so that it may be written in the form
@@
It is straightforward to verify, as shown in Appendix B, that $|\chi_{i}\rangle_{s}$ are the pointer states and that a generic state will evolve into a state whose system density matrix is diagonal in the pointer basis, i.e.
@@
(for a sufficiently large and scrambling environment) as is characteristic of einselection. It is also manifest from the form of Eq. (15) that $H$ is block diagonal in the pointer basis and, furthermore, that the $|\chi_{i}\rangle_{s}$ are robust to entanglement independent of the choice of the environment state or the size or scrambling nature of the environment. [^9] The distinction we are making here is that the robustness addresses the evolution of a state where the system is initially in the pure state $|\chi_{i}\rangle_{s}$, and Eq. (16) reflects the behavior of an arbitrary initial state. This feature can be seen in Fig. 2 where we show that randomizing the environment state does not degrade the purity of the reduced density matrix. We also note that the pointer states are stationary, while an arbitrary environment state will evolve under the evolution generated by the $H_{e}^{(i)}$ operators. All these properties, as well as a general recipe for finding this factorization for an arbitrary Hamiltonian, are proved in Appendix B.
We note that such a mapping to a factorization where the energy eigenstates are separable has some similarity with the case of dynamical decoupling in superconducting qubits (see e.g. Appendix A in [pokharel2018demonstration-lidar] for an overview). There, the “logical” qubits (what we call the “destination” TPS) correspond to the energy eigenstates of a world described by interacting “physical” qubits (what we have called the “fiducial” TPS). Defining the logical qubits in this way can mitigate certain sources of noise[heunisch2023tunable-hartmann].
In the decoherence literature authors often consider special Hamiltonians; such as those for which the system self-interaction commutes with the interaction, $H_{s}\otimes 1_{e},H_{\rm int}]=0$ or a “quantum measurement limit” where the Hamiltonian is dominated by a single interaction term $H\approx H_{s}\otimes H_{e}$. These are both special cases of the more general block-diagonalization procedure outlined above where the $H_{e}^{(j)}$ operators in Eq. (15) have specific relationships. For example, in the quantum measurement limit all the $H_{e}^{(i)}$ operators are the same up to an overall scaling, $H_{e}^{(j)}\propto H_{e}^{(i)}$. A particularly interesting sub-case of the block diagonal Hamiltonian is when there exists a decoherence-free subspace (DFS) (a topic relevant for quantum computing [Lidar_2003, quiroz2024dynamically-lidar-DFS]). In this factorization some, but not all, eigenvectors have vanishing entropy and, as in the quantum measurement limit, their corresponding $H_{e}^{(i)}$ operators are proportional to each other. Writing out the Hamiltonian in the global eigenbasis, we can interpret it as a sum of a decoupled (i.e. DFS) and an entangling part:
@@
Such a DFS solution lies in the intermediate regime of the general block diagonal form in Eq. (15) and the decoupled Hamiltonian in Eq. (14). Since $H_{\rm DFS}$ is subject to the same constraints as those for the (fully) decoupled Hamiltonian (see footnote 8), we do not expect to find such a factorization for random Hamiltonians. However, the central spin model admits both the decoupled and the block diagonal destinations, and in that case our algorithm is indeed able to find this intermediate DFS solution.
@@
This class of solutions is particularly interesting because it exhibits dynamic pointer states, though it fails to provide a complete basis of such states. Here, there is a dominant contribution to $|\psi\rangle_{w}$ from a low entropy energy eigenstate and smaller but, crucially, non-negligible contributions from high entropy eigenstates. More formally, in the language of Appendix B, the world energy eigenstates are of the form,
@@
where $|1\rangle_{w}$ has zero entropy, and we take the Schmidt coefficients $\beta_{j}^{(i)}$ to be generally nonzero for $j>1$ $\forall i\in\{2,..,d_{w}\}$, so the other energy eigenstates have nonzero entropy. (for convenience we choose to label the low entropy eigenstate $|1\rangle_{w}$). And the global state is of the form,
@@
This decomposition of $|\psi\rangle_{w}$ into eigenstates $|i\rangle_{w}$ and their corresponding linear entropy of entanglement, Eq. (8), is shown in Fig. 3 for a world with one system and two environment qubits  [^10] We note that the correlation between the entropy and $|\alpha_{i}|^{2}$ for $i>2$ seems typical of these ‘approximate eigenstate’ solutions but that it is likely a feature of the search algorithm and not of inherent physical significance. We have checked numerically that, as long as the condition $\alpha_{1}\gg\alpha_{j>1}$ is satisfied, there is considerable freedom in distributing the weights amongst the other higher entropy eigenstates.. In Fig. 2 c. we show the evolution of the purity of the reduced system density matrix of such a state (labeled ‘training $|\psi\rangle_{w}$’) compared to other states.
Figure 3: The decomposition of $|\psi\rangle_{w}$ in the energy eigenbasis for the case of the ‘approximate eigenstate” solutions discussed in Sec. IV.3 (see Eq. (20)) is shown in blue along with their corresponding linear entropy of each energy eigenstate (defined in Eq. (8)). These solutions are dominated by a single low entropy energy eigenstate, although high entropy eigenstates contribute to a small degree. The plot is arranged in order of decreasing $|\alpha_{i}|$.
The fact that the contribution to $|\psi\rangle_{w}$ from the high entropy eigenstates is small, but not negligible, is crucial as it leads to a pure $\rho_{s}$ for times several orders of magnitude greater than the decoherence time (see Fig. 2 c.), while allowing the state to be dynamic. These two behaviors resonate with our classical experience of subsystems, and cannot be achieved by the single low entropy eigenstate on its own.
It appears, at first, that such solutions are exactly the ones that the author of [tegmark2015consciousness] is in search of, i.e. those that have a high degree of predictability (as evidenced by the purity of the system state) as well as “autonomy” (that their dynamics are not trivial). We note that while this class of solutions satisfactorily minimizes our cost function (Eq. (12)), they do not yield a multi-dimensional distinguishable basis of pointer states of the form in Eq. (20). This should be contrasted with the block diagonal TPS (c.f. part (b)) which has a complete set of pointer basis states so that the system density matrix follows Eq. (16). However, in that case, the pointer states have trivial dynamics. While requiring a complete pointer basis seems important for standard discussions of einselection, singling out a special low entropy state (i.e. Eq. (19)) might be relevant for discussions of cosmological initial conditions.